A jet travels 2064 miles against a jetstream in 3 hours and 2514 miles with the jetstream in the same amount of time. What is the rate of the jet in still air and what is the rate of the jetstream?

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39631)

too common
ones like it been done a bunch of times on the site

Answer by ikleyn(52932)

.
A jet travels 2064 miles against a jetstream in 3 hours and 2514 miles with the jetstream
in the same amount of time. What is the rate of thae jet in still air and what is the rate of the jetstream?
~~~~~~~~~~~~~~~

Let u be the jet rate in still air (in miles per hour), and
let u be the jetstream rate.


Then the jet's effective speed flying with the jetstream is u+v miles per hour,
while its effective speed flying against the jetstream is u-v miles per hour.


From the condition, we have this system of equstions


    u + v = 2514/3 = 838  mph    (1)

    u - v = 2064/3 = 688  mph    (2)


To solve equations, add them. You will get


    2u = 838 + 688 = 1526;  hence  u = 1526/2 = 763  mph.


Then from equation (1),

    v = 838 - u = 838 - 763 = 75 mph.


ANSWER.  The jet rate in still air is 763 miles per hour;  the jeststream rate is 75 miles per hour.

Solved.

-------------

It is a typical "tailwind and headwind" word problem.

See the lessons
    - Wind and Current problems
    - Wind and Current problems solvable by quadratic equations
    - Selected problems from the archive on a plane flying with and against the wind
in this site, where you will find other similar solved problems with detailed explanations.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the section "Word problems",  the topic "Travel and Distance problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.